OpenStudy (anonymous):

The temperature in Buffalo, NY went from 0° to 17.3° in a 75-minute period. If the temperature increase was constant, how many degrees did the temperature rise per minute? Round your answer to the nearest hundredth. A.2.30° per minute B.0.23° per minute *MY CHOICE* C.4.34° per minute D.0.43° per minute

OpenStudy (anonymous):

\[\frac{ 17.3 }{ 75 } = \frac{ x }{ 1 }\] We should just have to cross multiply this proportion.

OpenStudy (anonymous):

@Pawtpie but wouldn't it be choice B

OpenStudy (anonymous):

I got: \[\frac{ 17.3 }{ 75 } = \frac{ x }{ 1 }\] \[17.3 = 75x\] \[{17.3 = 75x}\over75\] \[0.23 = x\]

OpenStudy (anonymous):

@Pawtpie Thank you, thats what me and my bro got too. just checking thanks again